OpenStudy (anonymous):

HELP PLEASE:) FAN&MEDAL&TESTIMONIAL:):) A committee of 5 people is to be chosen from a group of 8 women and 10 men. The committee must feature 3 women and two men? The committee must have more women than men?

OpenStudy (anonymous):

Without restrictions all the possibilities are accounted for by 18P5

OpenStudy (anonymous):

but with the above restrictions.....:(

ganeshie8 (ganeshie8):

Wrong, does the order in which you pick people matter here ?

OpenStudy (anonymous):

I don't think so.

ganeshie8 (ganeshie8):

The order in which you select the people doesn't matter here because, end of the day, all you bother about is "who" are in the committee. Not in which order you picked them.

ganeshie8 (ganeshie8):

So this better be a combination problem, yes ?

OpenStudy (anonymous):

Yes.

ganeshie8 (ganeshie8):

Without restrictions all the possibilities are accounted for by 18\(\color{red}{C}\)5

OpenStudy (anonymous):

What's the difference between P and C? I am sorry I am not very advanced

ganeshie8 (ganeshie8):

Maybe consider a quick example : Suppose you went to juice shop and ask for a salad made up of "bananas, apple, grapes"

ganeshie8 (ganeshie8):

does the order in which you ask matter here ?

OpenStudy (anonymous):

No. You can go about picking any

OpenStudy (anonymous):

But what difference does it make if permutation changed to combination?

OpenStudy (anonymous):

Lack of order?

ganeshie8 (ganeshie8):

`give me a salad made up of "bananas, apple, grapes"` is no different from `give me a salad made up of "apple, bananas, grapes"` order doesn't matter here, so this is a combination problem.

OpenStudy (anonymous):

ok I get it now.

ganeshie8 (ganeshie8):

consider another quick example : you want to choose four digits to set a password to your laptop hard disk : "8 7 3 9" does the order matter here ?

ganeshie8 (ganeshie8):

sorry to persist, but i think this is really important..

OpenStudy (anonymous):

No

ganeshie8 (ganeshie8):

Are you saying "8 7 3 9" is no different from "7 3 8 9" ?

ganeshie8 (ganeshie8):

does your computer accept/treat both passwords as same ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

The computer only accepts a given password you made at the boot up

OpenStudy (anonymous):

So the order would have to remain the same.

ganeshie8 (ganeshie8):

Exactly, since the "order" in which you enter the digits matter here, this is clearly a permutation problem.

ganeshie8 (ganeshie8):

That is how you differentiate between a permutation and combination problem

OpenStudy (anonymous):

Oh ok. I get it now.

OpenStudy (anonymous):

Ordinal or nominal

ganeshie8 (ganeshie8):

``` A committee of 5 people is to be chosen from a group of 8 women and 10 men. The committee must feature 3 women and two men? ``` How many ways can you choose 3 women from 8 women ?

OpenStudy (anonymous):

24

ganeshie8 (ganeshie8):

good guess, but no.

OpenStudy (anonymous):

Either this is too hard or I am too stupid:(

OpenStudy (anonymous):

2.7 something

OpenStudy (anonymous):

8/3=2.66

ganeshie8 (ganeshie8):

Neither, you just don't happen to have enough practice. You can think of this exactly same as the permutation problem. 3 women can be chosen from 8 women in 8C3 ways

ganeshie8 (ganeshie8):

where \(nCr\) is defined as \[nCr = \dfrac{n!}{r!*(n-r)!}\]

OpenStudy (anonymous):

8C3=number of possibilities where 3 women are picked from 8 women.

ganeshie8 (ganeshie8):

\[8C3 = \dfrac{8!}{3!*(8-3)!} = \dfrac{8!}{3!*5!}=\dfrac{8\times 7\times 6\times 5\times 4\times 3\times 2\times 1}{3\times 2\times 1*5\times 4\times 3\times 2\times 1}\]

ganeshie8 (ganeshie8):

Yes, see if you can simplify above

OpenStudy (anonymous):

(8*7*6*5)/3*2*1=336 1680/6=

OpenStudy (anonymous):

Because most numbers cancel each other out.

OpenStudy (anonymous):

8*7*6/3*2*1

ganeshie8 (ganeshie8):

Yes simplifies further

OpenStudy (anonymous):

330/6=55

ganeshie8 (ganeshie8):

8*7*6/3*2*1 8*7 56 right ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so 56 ways in which 3 women can be picked from 8 women.

ganeshie8 (ganeshie8):

Yes next find how many ways 2 men can be picked from the available 10 men

OpenStudy (anonymous):

2C10 so

ganeshie8 (ganeshie8):

10C2 ok

ganeshie8 (ganeshie8):

work it out

OpenStudy (anonymous):

10!/2!(10-2)!

OpenStudy (anonymous):

10!/2!*8!

OpenStudy (anonymous):

10*9*8*7*6*5*4*3*2*1/2*1*8*7*6*5*4*3*2*1*

OpenStudy (anonymous):

I simplify that

ganeshie8 (ganeshie8):

yes do it

OpenStudy (anonymous):

90/2=45

OpenStudy (anonymous):

45 ways in which 2 men can be picked from 10 men

OpenStudy (anonymous):

56*45 is my guess.

ganeshie8 (ganeshie8):

Excellent! so a committee with 3 women and two men can be formed in : \[8C3*10C2 = 56*45~\text{ways}\]

OpenStudy (anonymous):

I got it:)

OpenStudy (anonymous):

Do I use the same for the last one?

OpenStudy (anonymous):

where there are more men than women.

OpenStudy (anonymous):

there are 3 cases where men can be more than women 3 men 2 women 4 men 1 woman 5 men 0 women

OpenStudy (anonymous):

Sorry reverse of that

OpenStudy (anonymous):

more women than men. So do I multiply our previous answer by 3?

ganeshie8 (ganeshie8):

sry was away.. still here @Robert136

ganeshie8 (ganeshie8):

``` there are 3 cases where men can be more than women 3 women 2 men 4 women 1 man 5 women 0 man ``` Yes find the number of ways for each case, then add up.

ganeshie8 (ganeshie8):

you should get \[8C3*10C2 + 8C4*10C1+5C1*10C0\]